Find the inverse of f(x)=5x+1/6. Then, use your knowledge of composite functions to show that they are indeed inverses of each other. Show work
Answer by stanbon(75887)

Find the inverse of f(x)=5x+1/6. Then, use your knowledge of composite functions to show that they are indeed inverses of each other. Show work
----
Inverse: g(x) = (x-(1/6))/5
----
Proof:
f[g(x)] = f[(x-(1/6))/5] = 5[(x-(1/6))/5 + (1/6) = x-(1/6)+(1/6) = x
---------------
g[f(x)] = g[5x+(1/6)] = [5x+(1/6)-(1/6)]/5 = 5x/5 = x
=======================
Cheers,
Stan H.
=================